OpenStudy (anonymous):

Is the point (-4,5) a solution to the following system? 2x-15y=33 -3x+4y=-5

OpenStudy (anonymous):

To find out if it is or not, Sub the point \(P(-4,5)\) in for the first equation. If they both return true, then the point is a solution.

OpenStudy (anonymous):

2(-4)-15(5)=33

OpenStudy (anonymous):

So lets say: \(2x-15y=33\) \(2(-4)-15(5)=33\) \(-8-75=33\) \(-83\neq33\)

OpenStudy (anonymous):

I don't think it is true?

OpenStudy (anonymous):

You would be correct :-)

OpenStudy (anonymous):

that is what I got. Ok thank you for the much needed help!!

OpenStudy (anonymous):

Anytime! :-) @Algebra2013